# MaNGA DynPop – VII. A Unified Bulge-Disk-Halo Model for Explaining Diversity in Circular Velocity Curves of 6000 Spiral and Early-Type Galaxies

### v1

@@ Abstract
We derive circular velocity curves (CVCs) from stellar dynamical models for $\sim 6{,}000$ nearby galaxies in the final data release of the SDSS-IV MaNGA survey with integral-field spectroscopy, exploring connections between the inner gravitational potential (traced by CVC amplitude/shape) and galaxy properties. The maximum circular velocity ($V_{\rm circ}^{\rm max}$) and circular velocity at the half-light radius ($V_{\rm circ}(R_{\rm e}^{\rm maj})$) both scale linearly with the stellar second velocity moment $\sigma_{\rm e}^{2}\equiv\langle V^{2}+\sigma^{2}\rangle$ within the half-light isophote, following $V_{\rm circ}^{\rm max}\approx 1.75\sigma_{\rm e}$ (8% error) and $V_{\rm circ}(R_{\rm e}^{\rm maj})\approx 1.65\sigma_{\rm e}$ (9% error). CVC shapes (rising, flat, declining) correlate strongly with structural and stellar population properties: declining curves dominate in massive, early-type, bulge-dominated galaxies with old, metal-rich stars and early quenching, while rising CVCs prevail in disk-dominated systems with younger stellar populations and ongoing star formation. Using a unified bulge-disk-halo model, we predict CVC shapes with minimal bias, identifying three governing parameters: bulge-to-total mass ratio ($B/T$), dark matter fraction within $R_{\rm e}$, and bulge Sersic index. The distribution of CVC shapes across the mass-size plane reflects evolutionary pathways driven by (i) in situ star formation (spurring bulge growth) and (ii) dry mergers. This establishes CVC morphology as a diagnostic for galaxy evolution, linking dynamical signatures to structural and stellar population histories.
@@ 1 Introduction
The rotation curve (RC) of a galaxy represents the rotational velocity of stars or gas as a function of radial distance from the galaxy’s center. The first indirect evidence of dark matter (DM) came from studies of the velocity dispersion of galaxies within the Coma Cluster (Zwicky, 1933, 1937, 2009). Subsequent investigations into DM using galaxy RCs began with nearby galaxies. Babcock (1939) proposed that the nearly flat RC observed at the outskirts of M31 could be explained by an increasing mass-to-light ratio in the outer regions, while Oort (1940) identified substantial amounts of invisible matter in the outer parts of NGC 3115 based on its RC. With the improvements of observational instruments, the rotation curve (RC) has long been a powerful tool for studying the mass distributions of spiral galaxies, either using optical emission lines (Rubin & Ford, 1970; Rubin et al., 1980, 1982, 1985) or radio 21-cm line (Bosma & van der Kruit, 1979; Bosma, 1981).
@@
As indicators of the gravitational potential (encompassing both baryonic and DM mass distributions) within which galaxies reside, the amplitude and the shape of RCs are expected to correlate with various galaxy properties such as luminosity, morphology, and structure. For instance, the Tully-Fisher relation (Tully & Fisher, 1977) tightly links the maximum rotation velocities (the amplitudes) and the total luminosities (or stellar masses) of spiral galaxies. Beyond this well-known correlation between RC amplitude and mass, the shapes of RCs have also been found to be dependent on galaxy morphologies (Corradi & Capaccioli, 1990; Erroz-Ferrer et al., 2016), luminosities (Persic et al., 1996; Sofue & Rubin, 2001), or a combination of both (Casertano & van Gorkom, 1991; Noordermeer et al., 2007; Swaters et al., 2009; Yoon et al., 2021). These evident correlations imply that the gravitational potential plays a crucial role in the evolution of galaxy structure and morphology.
@@
In addition, the gravitational potential also has a significant impact on star-formation quenching, leaving imprints on the present-day stellar population properties. Various mechanisms that halt star formation of galaxies can be categorized into external (outside the galaxies) and internal (within the galaxies). The external mechanisms include ram pressure stripping (Gunn & Gott, 1972), strangulation (Larson et al., 1980), tidal stripping (Wetzel et al., 2013), dynamical heating from galaxy harassment (Moore et al., 1996), while internal mechanisms include but are not limited to halo quenching (Dekel & Birnboim, 2006), active galactic nucleus (AGN) feedback (Harrison, 2017), stellar feedback (Colling et al., 2018), mass quenching (Peng et al., 2010), bar quenching (Khoperskov et al., 2018), morphological quenching (Martig et al., 2009), and angular momentum quenching (Lu et al., 2022). Among these quenching mechanisms, the morphological quenching (Martig et al., 2009), also known as gravitational quenching (Genzel et al., 2014) or dynamical suppression (Davis et al., 2014; Gensior et al., 2020; Gensior & Kruijssen, 2021), occurs when the growth of a stellar spheroid or bulge stabilizes the gas disk instability (Toomre, 1964), thereby quenching star formation. This gravitational potential-related mechanism does not require gas removal (e.g. the tidal/ram pressure stripping) or massive halo (e.g. AGN feedback, shock heating induced halo quenching), and thus can explain the appearance of gas-rich but quenched galaxies in halos less massive than $10^{12}{\rm M_{\odot}}$. Linking the gravitational potential and stellar population properties to distinguish gravitational quenching from other quenching mechanisms is crucial in understanding the complex star-formation histories that shapes the evolution of galaxies.
@@
The RCs were initially derived from long-slit spectroscopy and only provided one-dimensional (1D) information on stellar kinematics or ionized gas kinematics. Due to random or systematic misalignment of the slit (i.e. if the slit is not aligned with the galaxy’s major axis), there might be an offset in the measured RCs, even if the galaxy is perfectly axisymmetric. The advent of integral field unit (IFU) surveys, e.g. SAURON (de Zeeuw et al., 2002), ATLAS3D (Cappellari et al., 2011), CALIFA (Sánchez et al., 2012), SAMI (Bryant et al., 2015), and MaNGA (Bundy et al., 2015), provides two-dimensional (2D) kinematic information, reducing the effect of slit misalignment when extracting RCs (e.g. Yoon et al., 2021; Ristea et al., 2024a).
@@
However, it was also pointed out that the RCs only work well in spirals and will underestimate the true CVCs of dispersion-dominated galaxies due to their significant non-circular motions (Roper et al., 2023; Downing & Oman, 2023; Sands et al., 2024). One way to overcome this intrinsic limitation of RCs and extend the studies to include early-type galaxies (ETGs) is applying corrections to account for the contribution of disordered motions, e.g. the asymmetric drift correction (Shetty et al., 2020; Bershady et al., 2024) or computing circular velocities from kinetic energy (Ristea et al., 2024b). A more direct and more accurate way is to constrain the gravitational potential (or CVCs) through detailed stellar dynamical models combined with spatially resolved stellar kinematics (e.g., Cappellari et al., 2013a; Li et al., 2017; Leung et al., 2018; Zhu et al., 2023b). CVCs derived from stellar dynamical models also have advantages over directly measured RCs, such as accounting for seeing effects and using accurately recovered inclination angle (Cappellari, 2008).
@@
In this paper, which is the seventh paper of our DynPop series, we derive the CVCs from the JAM models for 10000 galaxies (of which 6000 are deemed reliable) provided in the MaNGA DynPop project (Paper I; Zhu et al., 2023b). The MaNGA DynPop project not only provides the mass distributions but also the stellar population properties (Paper II; Lu et al., 2023), enabling us to study the dynamical scaling relations in conjunction with stellar population properties (Paper III; Zhu et al., 2024), the density profiles for galaxy groups and clusters by combining stellar dynamics and weak lensing (Paper IV; Wang et al., 2024), the variation of stellar initial mass function (IMF) in ETGs (Paper V; Lu et al., 2024), and a detailed comparison of total density slopes between MaNGA and simulations (Paper VI; Li et al., 2024). The structure of this paper is organized as follows. Section 2 briefly introduces the MaNGA data and how we derive the CVCs and other galaxy properties. In Section 3, we study the correlations between CVCs (including the amplitude and the shape) and other galaxy properties (Section 3.1 and Section 3.2), propose a model to quantify the shape of CVCs (Section 3.3), and investigate the evolution of CVC shapes on the mass-size plane (Section 3.4). We summarize our main results in Section 4. Throughout the paper, we assume a flat Universe with $\Omega_{\rm m}=0.307$ and $H_{0}=67.7\,\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$ (Planck Collaboration et al., 2016), for consistency with the other papers of the DynPop series.
@@ 2.1 The MaNGA survey
As one of the three projects in Sloan Digital Sky Survey-IV (SDSS-IV; Blanton et al., 2017), the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA) survey (Bundy et al., 2015) provides spatially resolved spectral measurements of $\sim 10000$ nearby galaxies. The MaNGA project uses the IFU technique to obtain spectra simultaneously across the face of target galaxies, employing tightly-packed fiber bundles that feed into the BOSS spectrographs (Smee et al., 2013; Drory et al., 2015) on the Sloan 2.5m telescope (Gunn et al., 2006). The field-of-view (FoV) of MaNGA observations extends radially out to 1.5 effective radii ($R_{\rm e}$) for about two-thirds of the galaxies (Primary+ sample), and up to 2.5 $R_{\rm e}$ for roughly one-third of the galaxies (Secondary sample) at higher redshifts (Law et al., 2015; Wake et al., 2017). The final sample exhibits an approximately flat stellar mass distribution across the range $10^{9}-6\times 10^{11}\,\rm M_{\odot}$ (Wake et al., 2017), with a median redshift of $z\sim 0.03$.
The MaNGA spectra cover a wavelength range of $3600-10300\,\rm{\AA}$, with a spectral resolution of $\sigma=72\,\rm{\rm km\,s^{-1}}$ (Law et al., 2016). The raw data are spectrophotometrically calibrated (Yan et al., 2016) and processed using the Data Reduction Pipeline (DRP; Law et al., 2016) to produce data cubes. Stellar kinematic maps are then extracted from these data cubes through the Data Analysis Pipeline (DAP; Belfiore et al., 2019; Westfall et al., 2019), which utilizes the ppxf software (Cappellari & Emsellem, 2004; Cappellari, 2017, 2023) and a subset of the MILES stellar library (Sánchez-Blázquez et al., 2006; Falcón-Barroso et al., 2011), MILES-HC, to fit the absorption lines in the IFU spectra. Before extracting stellar kinematics, the spectra are Voronoi binned (Cappellari & Copin, 2003) to a signal-to-noise ratio of $\rm S/N=10$ to obtain reliable measurements.
@@ 2.2 Dynamical models, CVCs, and other dynamical properties
We performed Jeans Anisotropic Modelling (JAM; Cappellari, 2008, 2020) to construct dynamical models for the whole MaNGA sample in Zhu et al. (2023b). The JAM model allows for anisotropy in second velocity moments and two different assumptions on the orientation of velocity ellipsoid, i.e. $\rm JAM_{cyl}$ (cylindrically-aligned) and $\rm JAM_{sph}$ (spherically-aligned). Four different mass models are adopted in Zhu et al. (2023b), which differ primarily in their assumptions about dark matter distributions: (i) the mass-follows-light model which assumes that the total mass density traces the luminosity density (hereafter MFL model), (ii) the model which assumes a spherical NFW (Navarro et al., 1996) dark halo (hereafter NFW model), (iii) the fixed NFW model which assumes a spherical NFW halo predicted by the stellar mass-to-halo mass relation in Moster et al. (2013) and mass-concentration relation in Dutton & Macciò (2014) (hereafter fixed NFW model), (iv) the model which assumes a generalized NFW (Wyithe et al., 2001) dark halo (hereafter gNFW model). Further details about the eight models can be found in Zhu et al. (2023b), while the data catalog is available online. [^1] https://manga-dynpop.github.io/.
Based on comparisons of observed and modeled stellar kinematics, the entire sample is classified according to different modeling qualities (Qual = $-1$, 0, 1, 2, 3 from worst to best). In this work, we select 6065 galaxies that are flagged as $\rm Qual\geqslant 1$, for which the dynamical quantities related to the total mass distribution are nearly insensitive to variations in model assumptions (Zhu et al., 2023b). Throughout this paper, we adopt the gNFW model, which is the most flexible mass model, with the $\rm JAM_{cyl}$ assumption (the results and conclusions remain consistent when adopting the $\rm JAM_{sph}$ assumption) unless stated otherwise. The total mass distribution consists of three components: the nuclear supermassive black hole, the stellar mass distribution, and the dark matter mass distribution (i.e. a gNFW dark halo in this model). The black hole mass is estimated from $M_{\rm BH}-\sigma_{\rm c}$ relation (McConnell et al., 2011), where $\sigma_{\rm c}$ is computed as the mean stellar velocity dispersion within 1 FWHM of MaNGA PSF. For the stellar component, we use the Multi-Gaussian Expansion (MGE; Emsellem et al., 1994; Cappellari, 2002) method to fit SDSS r-band images and obtain the surface brightness. Then the surface brightness is deprojected to obtain the luminosity density of the kinematic tracer in the three-dimensional space. The gNFW profile is written as
@@
where $r_{s}$ is the characteristic radius, $\rho_{s}$ is the characteristic density, and $\gamma$ is the inner density slope. For $\gamma=-1$, this function reduces to the NFW profile.
With the gravitational potential $\Phi(R,z)$ expressed in the form of MGE, we use the mge_vcirc procedure (equation 45 in Cappellari, 2020) in the jampy package to calculate the circular velocity $V_{\rm circ}$ at a given galactocentric radius $R$ in the equatorial plane. Assuming that the galaxies are axisymmetric, we further derive the maps with constant values of $V_{\rm circ}$ at a given galactocentric radius $R=\sqrt{x_{\rm bin}^{2}+y_{\rm bin}^{2}/q^{2}}$, where ($x_{\rm bin}$, $y_{\rm bin}$) are the coordinates of Voronoi bins and $q\equiv b/a$ is the axial ratio of half-light elliptical isophote derived from MGE. In Figure 1, we present examples of CVCs, circular velocity maps, line-of-sight velocity maps, stellar age maps and stellar metallicity maps for galaxies with different CVC shapes (rising, flat, or declining in the outskirts). Table 1 presents the parameters of CVCs, including the circular velocity at half-light radius and the maximum circular velocity within kinematic data range.
@@
[Mpc] [arcsec] [arcsec] [arcsec] [arcsec] [arcsec] [$\rm km~{}s^{-1}$] [$\rm km~{}s^{-1}$] [$\rm km~{}s^{-1}$] [$\rm km~{}s^{-1}$] [$\rm M_{\odot}$]
@@
Note. — (1) plateifu: The name of plate ID + IFU design ID (e.g. 7443-1901); (2) mangaid: Unique MaNGA ID (e.g. 12-84620); (3) DA: Angular diameter distance which assumes a flat Universe with $\Omega_{\rm m}=0.307$ and $h=0.677$ (Planck Collaboration et al., 2016); (4) $R_{\rm e}$: Effective radius (projected circular half-light radius from MGE fitting, in SDSS r band); (5) $R_{\rm e}^{\rm maj}$: Major axis of elliptical half-light isophote from MGE fitting, in SDSS r band; (6) $\rm rFWHM\_IFU$: The PSF FWHM values of IFU observations, in SDSS r band; (7) $R(V_{\rm circ}^{\rm max})$: The galactocentric radius where has the maximum circular velocity; (8) rmax: The kinematic data range, which is defined as the largest radius of the Voronoi bins; (9) $V_{\rm circ}(R_{\rm e})$: Circular velocity at half-light radius in the equatorial plane; (10) $V_{\rm circ}^{\rm maj}$: Circular velocity at the major axis of half-light ellipse in the equatorial plane; (11) $V_{\rm circ}^{\rm max}$: The maximum circular velocity within kinematic data range; (12) $V_{\rm circ}({\rm rmax})$: The circular velocity at rmax. (13) $\lg\,M_{\rm BH}$: The mass of central super massive black hole adopted in our dynamical models, which is derived from scaling relation (see Section 2.2); (14) Qual: The quality of JAM models ($-1$ to 3 from worst to best), only galaxies with $\rm Qual\geqslant 1$ have reliable CVC measurements.

### v2

# MaNGA DynPop. VII. A Unified Bulge-Disk-Halo Model for Explaining Diversity in Circular Velocity Curves of 6000 Spiral and Early-Type Galaxies

@@ Abstract
We derive circular velocity curves (CVCs) from stellar dynamical models for $\sim 6000$ nearby galaxies in the final data release of the SDSS-IV MaNGA survey with integral-field spectroscopy, exploring connections between the inner gravitational potential (traced by CVC amplitude/shape) and galaxy properties. The maximum circular velocity ($V_{\rm circ}^{\rm max}$) and circular velocity at the half-light radius ($V_{\rm circ}(R_{\rm e}^{\rm maj})$) both scale linearly with the stellar second velocity moment $\sigma_{\rm e}^{2}\equiv\langle V^{2}+\sigma^{2}\rangle$ within the half-light isophote, following $V_{\rm circ}^{\rm max}\approx 1.72\sigma_{\rm e}$ (7% error) and $V_{\rm circ}(R_{\rm e}^{\rm maj})\approx 1.62\sigma_{\rm e}$ (7% error). CVC shapes (rising, flat, declining) correlate strongly with structural and stellar population properties: declining curves dominate in massive, early-type, bulge-dominated galaxies with old, metal-rich stars and early quenching, while rising CVCs prevail in disk-dominated systems with younger stellar populations and ongoing star formation. Using a unified bulge-disk-halo model, we predict CVC shapes with minimal bias, identifying three governing parameters: bulge-to-total mass ratio ($B/T$), dark matter fraction within $R_{\rm e}$, and bulge Sersic index. The distribution of CVC shapes across the mass-size plane reflects evolutionary pathways driven by (i) in situ star formation (spurring bulge growth) and (ii) dry mergers. This establishes CVC morphology as a diagnostic for galaxy evolution, linking dynamical signatures to structural and stellar population histories.
@@ 1 Introduction
The rotation curve (RC) of a galaxy represents the rotational velocity of stars or gas as a function of radial distance from the galaxy’s center. The first indirect evidence of dark matter (DM) came from studies of the velocity dispersion of galaxies within the Coma Cluster (Zwicky, 1933, 1937, 2009). Subsequent investigations into DM using galaxy RCs began with nearby galaxies. Babcock (1939) proposed that the nearly flat RC observed at the outskirts of M31 could be explained by an increasing mass-to-light ratio in the outer regions, while Oort (1940) identified substantial amounts of invisible matter in the outer parts of NGC 3115 based on its RC. With the improvements of observational instruments, the RC has long been a powerful tool for studying the mass distributions of spiral galaxies, either using optical emission lines (Rubin & Ford, 1970; Rubin et al., 1980, 1982, 1985) or radio 21 cm line (Bosma & van der Kruit, 1979; Bosma, 1981).
@@
As indicators of the gravitational potential (encompassing both baryonic and DM mass distributions) within which galaxies reside, the amplitude and the shape of RCs are expected to correlate with various galaxy properties such as luminosity, morphology, and structure. For instance, the Tully-Fisher relation (Tully & Fisher, 1977) tightly links the maximum rotation velocities (the amplitudes) and the total luminosities (or stellar masses) of spiral galaxies. Beyond this well-known correlation between RC amplitude and mass, the shapes of RCs have also been found to be dependent on galaxy morphologies (Corradi & Capaccioli, 1990; Erroz-Ferrer et al., 2016), luminosities (Persic et al., 1996; Sofue & Rubin, 2001), or a combination of both (Casertano & van Gorkom, 1991; Noordermeer et al., 2007; Swaters et al., 2009; Yoon et al., 2021; Jeong et al., 2025). These evident correlations imply that the gravitational potential plays a crucial role in the evolution of galaxy structure and morphology.
@@
In addition, the gravitational potential also has a significant impact on star-formation quenching, leaving imprints on the present-day stellar population properties. Various mechanisms that halt star formation of galaxies can be categorized into external (outside the galaxies) and internal (within the galaxies). The external mechanisms include ram pressure stripping (Gunn & Gott, 1972), strangulation (Larson et al., 1980), tidal stripping (Wetzel et al., 2013), dynamical heating from galaxy harassment (Moore et al., 1996), while internal mechanisms include but are not limited to halo quenching (Dekel & Birnboim, 2006), active galactic nucleus (AGN) feedback (Harrison, 2017), stellar feedback (Colling et al., 2018), mass quenching (Peng et al., 2010), bar quenching (Khoperskov et al., 2018), morphological quenching (Martig et al., 2009), and angular momentum quenching (Lu et al., 2022). Among these quenching mechanisms, the morphological quenching (Martig et al., 2009), also known as gravitational quenching (Genzel et al., 2014) or dynamical suppression (Davis et al., 2014; Gensior et al., 2020; Gensior & Kruijssen, 2021), occurs when the growth of a stellar spheroid or bulge stabilizes the gas disk instability (Toomre, 1964), thereby quenching star formation. This gravitational potential-related mechanism does not require gas removal (e.g. the tidal/ram pressure stripping) or massive halo (e.g. AGN feedback, shock heating induced halo quenching), and thus can explain the appearance of gas-rich but quenched galaxies in halos less massive than $10^{12}{\rm M_{\odot}}$. Linking the gravitational potential and stellar population properties to distinguish gravitational quenching from other quenching mechanisms is crucial in understanding the complex star-formation histories that shape the evolution of galaxies.
@@
The RCs were initially derived from long-slit spectroscopy and only provided one-dimensional information on stellar kinematics or ionized gas kinematics. Due to random or systematic misalignment of the slit (i.e. if the slit is not aligned with the galaxy’s major axis), there might be an offset in the measured RCs, even if the galaxy is perfectly axisymmetric. The advent of integral field unit (IFU) surveys, e.g. SAURON (de Zeeuw et al., 2002), ATLAS3D (Cappellari et al., 2011), CALIFA (Sánchez et al., 2012), SAMI (Bryant et al., 2015), and MaNGA (Bundy et al., 2015), provides two-dimensional (2D) kinematic information, reducing the effect of slit misalignment when extracting RCs (e.g. Yoon et al., 2021; Ristea et al., 2024a).
@@
However, it was also pointed out that the RCs only work well in spirals and will underestimate the true CVCs of dispersion-dominated galaxies due to their significant non-circular motions (Roper et al., 2023; Downing & Oman, 2023; Sands et al., 2024). One way to overcome this intrinsic limitation of RCs and extend the studies to include early-type galaxies (ETGs) is applying corrections to account for the contribution of disordered motions, e.g. the asymmetric drift correction (Shetty et al., 2020; Bershady et al., 2024) or computing circular velocities from kinetic energy (Ristea et al., 2024b). A more direct and more accurate way is to constrain the gravitational potential (or CVCs) through detailed stellar dynamical models combined with spatially resolved stellar kinematics (e.g., Cappellari et al., 2013a; Li et al., 2017; Leung et al., 2018; Zhu et al., 2023b). CVCs derived from stellar dynamical models also have advantages over directly measured RCs, such as accounting for seeing effects and using accurately recovered inclination angles (Cappellari, 2008).
@@
In this paper, which is the seventh paper of our DynPop series, we derive the CVCs from the Jeans anisotropic modeling (JAM) models for 10,000 galaxies (of which 6000 are deemed reliable) provided in the MaNGA DynPop project (Paper I; Zhu et al., 2023b). The MaNGA DynPop project not only provides the mass distributions but also the stellar population properties (Paper II; Lu et al., 2023), enabling us to study the dynamical scaling relations in conjunction with stellar population properties (Paper III; Zhu et al., 2024), the density profiles for galaxy groups and clusters by combining stellar dynamics and weak lensing (Paper IV; Wang et al., 2024), the variation of stellar initial mass function (IMF) in ETGs (Paper V; Lu et al., 2024), and a detailed comparison of total density slopes between MaNGA and simulations (Paper VI; Li et al., 2024). The structure of this paper is organized as follows. Section 2 briefly introduces the MaNGA data and how we derive the CVCs and other galaxy properties. In Section 3, we study the correlations between CVCs (including the amplitude and the shape) and other galaxy properties (Section 3.1 and Section 3.2), propose a model to quantify the shape of CVCs (Section 3.3), and investigate the evolution of CVC shapes on the mass-size plane (Section 3.4). We summarize our main results in Section 4. Throughout the paper, we assume a flat Universe with $\Omega_{\rm m}=0.307$ and $H_{0}=67.7\,\mathrm{km~s^{-1}~Mpc^{-1}}$ (Planck Collaboration et al., 2016), for consistency with the other papers of the DynPop series. In this paper, $\rm log_{10}$ is denoted by $\rm lg$ to follow the official ISO 80000-2:2019 notation.
@@ 2.1 The MaNGA survey
As one of the three projects in Sloan Digital Sky Survey-IV (SDSS-IV; Blanton et al., 2017), the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA) survey (Bundy et al., 2015) provides spatially resolved spectral measurements of $\sim 10,000$ nearby galaxies. The MaNGA project uses the IFU technique to obtain spectra simultaneously across the face of target galaxies, employing tightly packed fiber bundles that feed into the BOSS spectrographs (Smee et al., 2013; Drory et al., 2015) on the Sloan 2.5m telescope (Gunn et al., 2006). The field of view (FoV) of MaNGA observations extends radially out to 1.5 effective radii ($R_{\rm e}$) for about two-thirds of the galaxies (Primary+ sample), and up to 2.5 $R_{\rm e}$ for roughly one-third of the galaxies (Secondary sample) at higher redshifts (Law et al., 2015; Wake et al., 2017). The final sample exhibits an approximately flat stellar mass distribution across the range $10^{9}-6\times 10^{11}\,\rm M_{\odot}$ (Wake et al., 2017), with a median redshift of $z\sim 0.03$.
The MaNGA spectra cover a wavelength range of $3600-10300\,\rm{\AA }$, with a spectral resolution of $\sigma=72\,\rm{\rm km\,s^{-1}}$ (Law et al., 2016). The raw data are spectrophotometrically calibrated (Yan et al., 2016) and processed using the Data Reduction Pipeline (DRP; Law et al., 2016) to produce data cubes. Stellar kinematic maps are then extracted from these data cubes through the Data Analysis Pipeline (DAP; Belfiore et al., 2019; Westfall et al., 2019), which utilizes the ppxf software (Cappellari & Emsellem, 2004; Cappellari, 2017, 2023) and a subset of the MILES stellar library (Sánchez-Blázquez et al., 2006; Falcón-Barroso et al., 2011), MILES-HC, to fit the absorption lines in the IFU spectra. Using a hierarchical-clustering approach, the entire MILES stellar library (985 spectra) is classified into 49 different clusters having similar spectra and therefore similar stellar parameters (e.g. effective temperature, metallicity, and surface gravity). For each cluster, all the spectra are normalized to a mean of unity and averaged without weighting to construct a representative stellar template. After excluding templates with prominent emission lines or relatively low $\rm S/N$, the final 42 stellar templates in the MILES-HC library enable reliable stellar kinematics measurements while reducing execution time by a factor of 25 (see more details in section 5 of Westfall et al., 2019). Before extracting stellar kinematics, the spectra are Voronoi binned (Cappellari & Copin, 2003) to a signal-to-noise ratio of $\rm S/N=10$ to obtain reliable measurements.
@@ 2.2 Dynamical models, CVCs, and other dynamical properties
We performed JAM (Cappellari, 2008, 2020) to construct dynamical models for the whole MaNGA sample in Zhu et al. (2023b). The JAM model allows for anisotropy in second velocity moments and two different assumptions on the orientation of the velocity ellipsoid, i.e. $\rm JAM_{cyl}$ (cylindrically-aligned) and $\rm JAM_{sph}$ (spherically-aligned). Four different mass models are adopted in Zhu et al. (2023b), which differ primarily in their assumptions about DM distributions: (i) the mass-follows-light model which assumes that the total mass density traces the luminosity density (hereafter MFL model), (ii) the model which assumes a spherical Navarro-Frenk-White (NFW; Navarro et al., 1996) dark halo (hereafter NFW model), (iii) the fixed NFW model which assumes a spherical NFW halo predicted by the stellar mass-to-halo mass relation in Moster et al. (2013) and mass-concentration relation in Dutton & Macciò (2014) (hereafter fixed NFW model), and (iv) the model which assumes a generalized NFW (Wyithe et al., 2001) dark halo (hereafter gNFW model). Further details about the eight models can be found in Zhu et al. (2023b), while the data catalog of dynamical properties is available online. [^1] As supplementary files of Zhu et al. (2023b) on the journal website and on the website of MaNGA DynPop at https://manga-dynpop.github.io/
Based on comparisons of observed and modeled stellar kinematics, the entire sample is classified according to different modeling qualities (Qual = 1-1, 0, 1, 2, 3 from worst to best). In this work, we select 6065 galaxies that are flagged as Qual $\geqslant 1$, for which the dynamical quantities related to the total mass distribution are nearly insensitive to variations in model assumptions (Zhu et al., 2023b). Throughout this paper, we adopt the gNFW model, which is the most flexible mass model, with the $\rm JAM_{cyl}$ assumption (the results and conclusions remain consistent when adopting the $\rm JAM_{sph}$ assumption) unless stated otherwise.
[…]
@@
where $r_{s}$ is the characteristic radius, $\rho_{s}$ is the characteristic density, and $\gamma$ is the inner density slope. This profile follows the same functional form as the commonly used gNFW profile (e.g. equation 1 in Wyithe et al., 2001), while its amplitude linearly scales by a factor of $\left(\frac{1}{2}\right)^{-\gamma-3}$ for a given $\gamma$ ($\rho_{s}$ and $\gamma$ are independent parameters).
With the JAM-determined gravitational potential $\Phi(R,z)$ expressed in the form of MGE, we use the mge_vcirc procedure (equation 45 in Cappellari, 2020) in the jampy package to calculate the circular velocity $V_{\rm circ}$ at a given galactocentric radius $R$ in the equatorial plane. Assuming that the galaxies are axisymmetric, we further derive the maps with constant values of $V_{\rm circ}$ at a given galactocentric radius $R=\sqrt{x_{\rm bin}^{2}+y_{\rm bin}^{2}/q^{2}}$, where ($x_{\rm bin}$, $y_{\rm bin}$) are the coordinates of Voronoi bins and $q\equiv b/a$ is the axial ratio of a half-light elliptical isophote derived from MGE. In Figure 1, we present examples of CVCs, circular velocity maps, line-of-sight velocity maps, stellar age maps and stellar metallicity maps for galaxies with different CVC shapes (rising, flat, or declining in the outskirts). Table 1 presents the parameters of CVCs, including the circular velocity at half-light radius and the maximum circular velocity within kinematic data range. In Table 2, we provide the systematic uncertainties arising from model differences, following the calculation presented in table 3 of Zhu et al. (2023b).
@@
(Mpc) (arcsec) (arcsec) (arcsec) (arcsec) (arcsec) ($\rm km~s^{-1}$) ($\rm km~s^{-1}$) ($\rm km~s^{-1}$) ($\rm km~s^{-1}$) ($\rm M_{\odot}$)
@@
Note. — The dynamical quantities in this table are derived from $\rm JAM_{cyl}$ + gNFW model (see Section 2.2). (1) plateifu: The name of plate ID + IFU design ID (e.g. 7443-1901); (2) mangaid: unique MaNGA ID (e.g. 12-84620); (3) DA: angular diameter distance that assumes a flat Universe with $\Omega_{\rm m}=0.307$ and $h=0.677$ (Planck Collaboration et al., 2016); (4) $R_{\rm e}$: effective radius (projected circular half-light radius from MGE fitting, in SDSS $r$ band); (5) $R_{\rm e}^{\rm maj}$: major axis of elliptical half-light isophote from MGE fitting, in SDSS $r$ band; (6) $\rm rFWHM\_IFU$: the PSF FWHM values of IFU observations, in SDSS $r$ band; (7) $R(V_{\rm circ}^{\rm max})$: the galactocentric radius where CVC has the maximum circular velocity; (8) rmax: the kinematic data range, which is defined as the largest radius of the Voronoi bins; (9) $V_{\rm circ}(R_{\rm e})$: circular velocity at half-light radius in the equatorial plane; (10) $V_{\rm circ}^{\rm maj}$: circular velocity at the major axis of half-light ellipse in the equatorial plane; (11) $V_{\rm circ}^{\rm max}$: the maximum circular velocity within kinematic data range; (12) $V_{\rm circ}({\rm rmax})$: the circular velocity at rmax; (13) $\lg\,M_{\rm BH}$: the mass of central supermassive black hole adopted in our dynamical models, which is derived from scaling relation (see Section 2.2); (14) Qual: the quality of JAM models (1-1 to 3 from worst to best), only galaxies with Qual $\geqslant 1$ have reliable CVC measurements.